OpenStudy (peterg2001):

help please will medal check if this is right In order to receive credit, you must SHOW ALL YOUR WORK that is necessary to answer the question Question 1: Solve-3y+6_<_-15 and graph the solution on a number line. <-----------------------------------> Question 2: Evaluate 3/21 divid 6/7 and write in lowest terms. Question 3: Evaluate -3 5/8+7 3/4 and write as a mixed number

OpenStudy (peterg2001):

(2) (3/21)/(6/7) = invert and multiply: (3/21)(7/6) = 21/126 = 1/6 (3) -3 5/8 + 7 3/4 = improper fractions: -29/8 + 31/4 = common denominators: -29/8 + 62/8 = add: 33/8 = mixed number: 4 1/8

OpenStudy (anonymous):

Yep its riht

OpenStudy (peterg2001):

sorry missed first one @King_Trae_OKC 1) -3y+6<-15 -3y<-21 y>-21/-3 y>7 Then draw that on the number line- everything greater than 7.

OpenStudy (peterg2001):

@ThisIsAdriAtHerBest @Crissy15

OpenStudy (peterg2001):

@abram134 @Directrix

OpenStudy (peterg2001):

@geneticrockhopper247

OpenStudy (peterg2001):

@King_Trae_OKC

OpenStudy (anonymous):

Are you asking if 7 is the answer?

OpenStudy (peterg2001):

know i ask to check if thats all steps @King_Trae_OKC

OpenStudy (anonymous):

Yeah thats all the steps lmao

OpenStudy (peterg2001):

thanks

OpenStudy (anonymous):

No problem